OpenStudy (anonymous):

Please help!! Find any points of intersection between f(x)=2x+3 amd g(x)=4/(x) -1 Leave your answers in exact form. I've tried substitution, but think I'm not doing it right. Appreciate any help :)

OpenStudy (anonymous):

this is what i got when I tried to graph them.

OpenStudy (anonymous):

So just equate the two and solve for x, \[2x+3=4/x -1 => 2x^2+4x-4=0 => x^2+2x-2=0\] so they intersect at two points just like what the graph above is showing, all you need now is to solve for x using the quadratic formula and substitute both in one of the equations to get the ordinates of the points.

OpenStudy (anonymous):

oh yeah! forgot about using the quadratic formula! haha, i'll try that now @gagafacea1

OpenStudy (anonymous):

great!